OpenStudy (anonymous):

For x≥2, y≥4 and x+y≤16, what is the maximum value of y/x

OpenStudy (hoblos):

y/x is maximum when y is maximum and x is minimum x is minimum at x=2 y is maximum when x+y=16 so y=14 max value of y/x=14/2 = 7